# Privacy Preservation Techniques (PPTs) in IoT Systems: A Scoping Review and Future Directions Author’s draft for soliciting feedback - February 20, 2025

### v1

@@ IX-E RQ5: What privacy types are the main focus of most PPTs in IoT systems?
According to our review, the three main privacy types in IoT are user, device, and location privacy. Some studies proposed that PPTs can protect the privacy of both the user and the device, as seen in Table V. However, our observation shows that most privacy implementations in IoT devices mainly focus on protecting user data stored in these IoT devices, with very limited or no emphasis on the importance of preserving the identity of the IoT device that stores these user data. This observation shows that user privacy is the ultimate goal of many research studies on PPT in IoT systems.
IX-F Relationship between Key concepts in IoT system privacy
@@
In this scoping review, we proposed a privacy preservation taxonomy that includes major privacy entities, as explained in Section II. This taxonomy encompasses Privacy Enhancing Technologies (PET), Privacy Protection Techniques (PPT), Privacy Threats, Privacy Goals, and the technologies used to develop PET. Although previous studies have mentioned these entities, none have explored the correlations among them in the context of privacy preservation, particularly within an IoT environment. Figure 9 illustrates the relationships among these privacy entities and detailed examples for each entity.
X Future Direction
@@
After proper reviews of PPTs proposed in the selected papers from top literature databases, this study observed some future research areas that need researchers’ attention.
X-A Need for device privacy preservation (Device Identity protection)
@@
Finn et al. [11] and Liyanage et al. [12] have proposed several types of privacy. However, in a typical user service, provider interactions within an IoT system use two major types of information flow: (a) user data, such as personally identifiable information (PII), including contacts, names, photos, and age, and (b) IoT device data and activity, such as IP address, geolocation, and device ID. These two major data are essential; however, protecting the identity of the device is crucial, as its protection could directly impact the privacy of the user and the location [186].
@@
Research Gap Despite limited studies on device data privacy, most studies are centered around protecting user data stored, processed, or transmitted across devices in IoT. Therefore, research on device identity privacy is essential for protecting sensitive information about devices in IoT systems.
X-B Need for proper understanding of privacy threats activities
Providing PPT to mitigate privacy threats in IoT is important, but more important is understanding the activities of these privacy threats. The knowledge of privacy threat activities gives more insight to provide proper PPT for mitigating these privacy threats.
Research Gap Implementing a comprehensive PTA for IoT systems is essential to address the full spectrum of privacy threats. In addition, this study observed a lack of privacy-based datasets to train and test known privacy threats in the Internet of Things. As an interim solution, generating synthetic datasets could simulate typical threat behaviors in IoT systems, helping to build and validate effective privacy protection measures.
@@ X-B Need for proper understanding of privacy threats activities
X-C Importance of Privacy Engineering (PE), Privacy Threat Analysis (PTA), and Privacy Impact Assessment (PIA) in privacy preservation
PE is the continuous and overall implementation of privacy protection across any IoT system. PE is required in IoT systems to implement privacy protection beyond the design stage. PTA analyzes privacy threats, and PIA assesses the risk level of these privacy threats in IoT systems. These three privacy approaches (PE, PTA, and PIA) are essential in any IoT system. However, an integrated approach that incorporates these privacy tools is essential in selecting appropriate PPTs to mitigate privacy threats in IoT systems.
Research Gap PE, PTA, and PIA integrated approach is important at an early stage of implementing privacy preservation in IoT applications. However, researchers have not effectively emphasized such approach in implementing privacy protection in a wide range of IoT systems. Alalade et al. [305] implemented this integrated approach in the Smart Home (SH) environment. The end result of their work ranked privacy threats based on their level of impact in an SH, enabling suggestions of possible PPTs to mitigate some of these threats. Applying this and similar approaches to other IoT applications should enable proper and in-depth privacy protection.
X-D The influence of emerging technology
@@
The introduction of emerging technologies has a dual impact on IoT systems. On the one hand, technology plays a crucial role in improving privacy protection [6, 7]. On the other hand, it also provides malicious actors with advanced tools to carry out privacy threats [4, 5]. This duality highlights the complex balance between utilizing technological advancements for privacy protection and managing the increased risks presented by the sophisticated tools employed by malicious individuals.
@@
Research Gap Researchers should leverage emerging technology to enhance privacy protection in IoT systems. For instance, integrating an appropriate lightweight post-quantum cryptography algorithm with the assistance of generative AI will provide quantum-safe privacy in an IoT system.
XI Conclusion
@@
This scoping review utilizes PRISMA-ScR guidelines and Arksey et al.’s framework to review PPTs in IoT systems comprehensively.
@@
Out of 725 papers collected, 243 met the selection criteria. These selected papers were analyzed for insights into PPT goals, PET building technologies, PPT implementation in IoT applications, where these PPTs are implemented in IoT computing architectures, and protected privacy types.
@@
Our review discovered that cryptography is the predominant technology for preserving privacy in IoT systems. At the same time, anonymity remains the primary objective of PPT in IoT implementations, addressing our research questions RQ1 and RQ2, respectively. Furthermore, the continuous expansion of IoT applications across diverse domains, with new introductions in social networks as SIoT and multimedia networks as (IoMT) reflects the growing influence of IoT, which answers our RQ3 research question. This study also reveals that PPT is implemented across various IoT computing layers: cloud, fog, edge, and local network. However, most PPT implementation occurs at the cloud layer of the IoT system. In addition, medical IoT has more publications on the implementation of PPT than other IoT applications, as gathered in our study, addressing our research question RQ4. This study discovered that while most PPT implementations prioritize user privacy, fewer studies focus on device and location privacy, thus answering RQ5.
@@
Finally, implementing effective PPT to address privacy threats in IoT systems can be challenging due to IoT system resource constraints commonly associated with these systems. Nevertheless, this study highlights our findings through a comprehensive scoping review, outlines future research directions in PPT implementation in IoT systems, and acknowledges limitations in this work. This review aims to serve as a valuable foundation for researchers in selecting appropriate PPTs for IoT applications and provide an understanding of the privacy types these PPTs protect.
@@
Limitations : The study focused on articles published from 2010 to early 2023, potentially missing relevant studies outside this timeframe. The inclusion criteria limited articles to English-language publications, which may have excluded significant non-English research. Additionally, as a scoping review, the study provided a broad overview rather than an in-depth analysis, potentially omitting essential nuances. These limitations should be considered when interpreting the findings.

### v2

@@ 9.5 RQ5: What privacy types are the main focus of most PPTs in IoT systems?
According to our review, the three main privacy types in IoT are user, device, and location privacy. Some studies proposed that PPTs can protect the privacy of both the user and the device, as seen in Table 5. However, our observation shows that most privacy implementations in IoT devices primarily focus on protecting user data stored in these devices, with very limited or no emphasis on preserving the identity of the IoT device that stores this data. This observation indicates that user privacy is the primary objective of many research studies on PPT in IoT systems.
9.6 Relationship between Key concepts in IoT system privacy
@@
In this scoping review, we proposed a privacy preservation taxonomy that includes major privacy entities, as explained in Subsection 2.5. This taxonomy encompasses privacy-enhancing technologies (PET), privacy preservation techniques (PPT), Privacy Threats, Privacy Goals, and the technologies used to develop PET. Although previous studies have mentioned these entities, none have explored the correlations among them in the context of privacy preservation, particularly within an IoT environment. Figure 4 illustrates the relationships among these privacy entities and detailed examples for each entity.
10 Future Direction
@@
After reviewing the PPTs proposed in the selected papers from top literature databases, we identified some future research areas that require the attention of researchers.
10.1 Need for device privacy preservation (Device identity protection)
@@
Finn et al. [155] and Liyanage et al. [299] have proposed several types of privacy. However, in a typical user service, provider interactions within an IoT system use two major types of information flow: (a) user data, such as personally identifiable information (PII), including contacts, names, photos, and age, and (b) IoT device data and activity, such as IP address, geolocation, and device ID. These two major data are essential; however, protecting the identity of the device is crucial, as its protection could directly impact the privacy of the user and the location [21] [22]. It is essential to note that a compromised IoT device will holistically impact the user and device, as the IoT device itself stores, processes, and transmits these data.
@@
Research Gap Despite limited studies on device data privacy, most studies on device privacy still centered around protecting user data stored, processed, or transmitted across these devices in IoT. Therefore, research on device identity privacy is essential for protecting sensitive information about devices in IoT systems. We will be considering both the user and device data privacy in subsequent chapters in this thesis.
10.2 Need for proper understanding of privacy threats activities
While providing PPTs to mitigate IoT privacy threats is essential, understanding the underlying activities and mechanisms of these threats through systematic privacy threat analysis (PTA) is equally critical. Such analysis provides the foundational insights necessary to design and implement effective mitigation strategies that address the root causes of privacy vulnerabilities.
Research Gap
Our scoping review revealed a significant gap: most reviewed papers proposing PPTs did not conduct prior analysis of the specific privacy threats they aimed to mitigate. This analytical deficiency likely explains why many reviewed PPTs papers fail to specify the targeted computing layer or privacy type explicitly in their work. A comprehensive PTA systematically identifies vulnerable areas within IoT architectures where privacy threats are most likely to occur, thereby enabling targeted and effective PPT deployment in these critical locations.
Implementing comprehensive PTA for IoT systems is essential to address the full spectrum of privacy threats, particularly by characterizing threat actors and their behavioral patterns. Machine learning approaches offer a promising avenue for automating threat analysis by training models on privacy-preserving datasets. However, we identified a lack of privacy-focused datasets for training and testing against known privacy threats in IoT systems. To address this methodological gap, a multi-faceted approach can be proposed by generating a synthetic dataset that can simulate realistic threat behaviors in IoT environments, facilitating the development and validation of robust privacy protection measures. Alternatively, expert elicitation through structured user studies with privacy specialists can provide qualitative insights into emerging threats, attack tactics, and evasion techniques, thereby enriching our understanding of the evolving threat landscape.
@@ 10.2 Need for proper understanding of privacy threats activities
10.3 Importance of Privacy Engineering (PE), Privacy Threat Analysis (PTA), and Privacy Impact Assessment (PIA) in privacy preservation
Privacy measures such as PPT are typically implemented once, during design or during IoT system setup. However, privacy measures should extend beyond the design stage, as many vulnerabilities can arise when the system is operational and are exploitable by a privacy threat. Therefore, the privacy threat mitigation approach should be periodically reviewed, particularly when changes are made to the IoT system, thereby engineering the process of privacy preservation.
Privacy engineering (PE) represents the continuous and comprehensive implementation of privacy protection throughout the lifecycle of IoT systems, extending beyond the initial design phase to encompass operational deployment and maintenance. Within the PE framework proposed by [176], privacy threat analysis (PTA) systematically identifies and analyzes privacy threats, while privacy impact assessment (PIA) evaluates their associated risk levels. These two privacy approaches (PTA and PIA) are fundamental pillars of PE implementation in any operational system. However, an integrated approach that incorporates these privacy frameworks is essential not only for implementing proactive privacy measures but also for maintaining adaptive, up-to-date privacy protection in IoT systems.
Research Gap Integrating privacy measures is crucial in the early stages of an IoT system; however, these measures should extend beyond the design stage. Privacy measures should be incorporated during installation and configuration, and maintained throughout the production use of the IoT system. However, researchers have not effectively emphasized how privacy measures can be implemented continuously beyond the design stage. Introduction of the PE approach, through PTA and PIA, would ensure that privacy is implemented at every stage of the IoT lifecycle, thereby keeping privacy preservation up to date in the IoT system. In the event of any change to the IoT system, whether by adding a new device or modifying configuration, the PTA and PIA should be repeated to ensure that privacy preservation remains up to date. The result will enable prioritization of privacy threats based on their potential risk in IoT, allowing recommendations for possible PPTs to mitigate these threats. Applying this to IoT applications such as Smart Home system should enable proactive and in-depth privacy protection.
10.4 The influence of emerging technology
@@
Evidence from our scoping review indicates that the introduction of emerging technologies, such as post-quantum cryptography, has contributed to privacy preservation in IoT systems by improving privacy protection [110, 323]. On the other hand, it also provides malicious actors with advanced tools to carry out privacy threats [184, 99]. This duality highlights the complex balance between utilizing technological advancements for privacy protection and managing the increased risks presented by the sophisticated tools employed by malicious individuals.
@@
Research Gap Researchers should leverage emerging technology such as post-quantum algorithms and Large Language Model (LLM) to enhance privacy protection in IoT systems. For instance, integrating an appropriate, lightweight post-quantum cryptography algorithm, with the assistance of generative AI, will provide quantum-safe privacy in an IoT system.
11 Conclusion
@@
This scoping review employs the PRISMA-ScR guidelines and Arksey et al.’s framework to comprehensively review PPTs in IoT systems.
@@
Out of 1177 papers collected, 329 met the selection criteria. These selected papers were analyzed for insights into PPT goals, PET building technologies, PPT implementation in IoT applications, where these PPTs are implemented in IoT computing architectures, and protected privacy types.
@@
Our review revealed that cryptography is the primary technology used to maintain privacy in IoT systems. At the same time, anonymity remains the primary objective of PPT in IoT implementations, addressing our research questions RQ1 and RQ2, respectively. Furthermore, the continuous expansion of IoT applications across diverse domains, with new introductions in social networks as SIoT and multimedia networks as IoMT, reflects the growing influence of IoT, which answers our RQ3 research question. This study also reveals that PPT is implemented across various IoT computing layers, including cloud, fog, edge, and local networks. However, most PPT implementation occurs at the cloud layer of the IoT system. In addition, medical IoT has more publications on the implementation of PPT than other IoT applications, as our study has gathered, addressing our research question RQ4. We discovered that while most PPT implementations prioritize user privacy, fewer studies focus on device and location privacy, thus answering RQ5.
@@
Finally, implementing effective PPT to address privacy threats in IoT systems can be challenging due to IoT system resource constraints commonly associated with these systems. Nevertheless, this study highlights our findings through a comprehensive scoping review, outlines future research directions for PPT implementation in IoT systems, and acknowledges the limitations of this work. This review aims to serve as a valuable foundation for researchers in selecting appropriate PPTs for IoT applications and to provide an understanding of the types of privacy these PPTs protect.
@@
Limitations: The study focused on articles published from 2010 to early 2025, potentially missing relevant studies outside this timeframe. The inclusion criteria limited articles to English-language publications, which may have excluded significant non-English research. Additionally, as a scoping review, the study provided a broad overview rather than an in-depth analysis, which may have potentially omitted essential nuances. These limitations should be considered when interpreting the findings.
Acknowledgement
This work was partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) through the NSERC Discovery Grant program.